About a special class of quasigroups H. Zeitler, Bayreuth, Germany

Keywords: quasigroups, Steiner–systems of block size 5 AMS Subject classification: 05B05 Abstract: In the axiomatic approach to the investigation of Steiner–systems, e.g. S(2, 4, v), quasigroups can be important tools. In this paper systems S(2, 5, v) are treated in a similar way.